# Signed Networks: theory, methods, and applications

### v1

@@ 6.2.1 In-person social communities
  • Sampson’s monastery data [sampsonNovitiatePeriodChange1968] records the evolution of relations between 18 monks during a political crisis within a US cloister, across three temporal waves. Each wave, the monks were asked to select the three others they esteemed the most and the one they esteemed the least. Similar ratings were collected for liking, influence, and praise instead of esteem.
  • Nordlie-Newcomb’s fraternity dataset [nordlieLONGITUDINALSTUDYINTERPERSONAL1958, newcombAcquaintanceProcess1961] is the result of an experiment of social psychology conducted at the University of Michigan. Seventeen male college students were housed together, and the authors periodically collected data via surveys to assess the evolution of friendships and antagonisms between the students. A similar study collected data about college female students in the US [Lemann1952].
  • Spanish high school students [ruiz-garciaTriadicInfluenceProxy2023]. Recently, Ruiz-Garcia et al. [ruiz-garciaTriadicInfluenceProxy2023] surveyed more than three thousand students (11-15 years old) in 13 Spanish schools, asking them about which other students they liked or disliked. The result is a directed, weighted network with over 60,000 edges, with additional information about individual features and personality traits.
  • Behavioural observations on preschoolers [nealInferringSignedNetworks2022]. Neal et al. [nealInferringSignedNetworks2022] inferred positive and negative links between children in two classrooms of preschoolers ($N=35$), based on behavioural patterns during playtime. To do so, they leverage the backbone procedure described in [nealBackboneBipartiteProjections2014] (see Section 6.1.2).
Other relevant data sources include signed networks of relationships between antique Philosophers [Lemann1952], alliances and antagonisms between $N=16$ tribes of the Highlands region in New Guinea [readCulturesCentralHighlands1954], as well as relationships between inhabitants of $N=32$ Honduran villages collected during a clinical trial. We finally mention another widely used source of data, the Urban Communes Dataset, which records the nature of relationships between groups of inhabitants in several major US cities, alongside socio-demographic indicators and individual beliefs, across multiple waves. Formerly freely available, it is now accessible on request.
@@ 6.2.2 Online platforms
The availability of large-scale, high-quality signed datasets for studying online social interactions remains limited. A few datasets have become standard benchmarks, in particular those hosted in the SNAP repository [snapnets], and have supported a wide range of work on sign prediction, community detection, and structural balance:
  • Epinions trust network [leskovecSignedNetworksSocial2010]. This dataset originates from the Epinions product review platform, where users could explicitly express trust or distrust toward other members. The resulting signed network captures these declared trust relationships among users.
  • Slashdot Zoo 313131www.slashdot.org [gomez2008, leskovecSignedNetworksSocial2010]. Slashdot is a technology-oriented news platform featuring user-submitted and editor-evaluated stories related to science and technology. In 2002, the site introduced the Zoo feature, allowing users to tag others as friends (those whose comments they appreciated) or foes (those whose comments they found uninteresting). The dataset includes three temporal snapshots of this signed social network.
Both Epinions [liuPredictingTrustsUsers2008, leskovecPredictingPositiveNegative2010, hamedaniTrustRecEffectiveApproach2021, ahmadalinezhadSignPredictionSigned2018, papaoikonomouPredictingEdgeSigns2014] and Slashdot [kunegisSlashdotZooMining2009, ahmadalinezhadSignPredictionSigned2018, papaoikonomouPredictingEdgeSigns2014] have been widely used in the literature for sign prediction and structural balance studies.
  • Bitcoin Alpha and Bitcoin OTC [kumarCommunityInteractionConflict2018]. These datasets represent who-trusts-whom networks extracted from two Bitcoin trading platforms. Users rate each other following transactions, creating directed weighted links that indicate trust or distrust. Since participants are anonymous, the reputation systems underlying these platforms aim to mitigate fraud and risk.
  • Wikipedia Requests for Adminship (RfA) [leskovecPredictingPositiveNegative2010, leskovecSignedNetworksSocial2010, west2014exploiting]. This dataset captures the voting process through which Wikipedia editors decide whether to promote a candidate to administrator status. Each vote can express support, opposition, or neutrality. The initial version, based on a 2008 dump, includes nearly 2,8002{,}800 elections. A later extended version by West et al. [west2014exploiting] spans the entire period from 2003 to 2013 and also contains the textual content of votes (i.e., the accompanying comments). Besides RfA, several studies have inferred signed relations from co-editing behaviour and interaction patterns among editors (see Section 6.1.4).
  • Reddit Hyperlink Network [kumarCommunityInteractionConflict2018]. This dataset represents interactions between subreddits over a 2.5-year period (2014–2017). It is constructed from hyperlinks in user posts that reference content in other subreddits. Each edge originates from a source subreddit and points to a target subreddit, with sentiment annotations derived from the text of the linking post, resulting in a signed inter-community network.
@@
  • DEBAGREEMENT [pougue-biyongDEBAGREEMENTCommentreplyDataset]. This dataset consists of 42,894 annotated comment–reply pairs collected from five politically and socially active subreddits: r/BlackLivesMatter, r/Brexit, r/climate, r/democrats, and r/Republican. Each pair is labelled as agree, neutral, or disagree, creating a signed interaction network among users based on the tone of their replies.
  • Birdwatch (now Community Notes) [yasseri2023can, allenBirdsFeatherDont2022]. This platform was launched by Twitter in January 2021 as a crowd-sourced fact-checking platform where selected users rated Tweet trustworthiness through evaluative notes. Signed networks have been extracted from user-note ratings from Birdwatch data to study political polarisation and antagonism [yasseri2023can, allenBirdsFeatherDont2022]. As of this writing, anonymised Birdwatch (now Community Notes) data remains publicly available.
  • DerStandard Forum Interactions [fraxanet2025decade, derstandard_data]. DerStandard  [^32] https://www.derstandard.at/ is an Austrian online newspaper that features a popular comment section integrated with its articles. The dataset captures user interactions on the platform over nearly ten years (2013–2022), including explicit agreement and disagreement votes on comments. These interactions can be used to build temporal signed networks across diverse topics such as football, national elections, or COVID-19 policy [fraxanetUnpackingPolarizationAntagonism2024].
  • Meneame.net. Meneame  [^33] https://www.meneame.net/, a Spanish news aggregator with a similar structure to Reddit, was studied by [candellone2025community, candellone2025negative] to investigate the role of negative interactions in highlighting polarised content. The platform allows users to submit news stories and interact through upvotes, downvotes, and comments. Moreover, unlike many platforms, Menéame applies no personalised recommendations, so any observed phenomenon arises purely from user behaviour.
Some additional platforms have supported signed or sign-inferable interactions, though the corresponding research does not provide public access to the datasets. Examples are from Yahoo Answers [yangFriendFrenemyPredicting2012], where the authors infer trust and distrust through question-and-answer behaviour and voting patterns; Essembly.com [brzozowskiFriendsFoesIdeological2008], where users can label others as “friends” or “foes”; Pardus [szellMultirelationalOrganizationLargescale2010], a multi-relational online game, included both cooperative and antagonistic links.
@@ 6.2.3 Political systems
The formalism of signed networks has raised interest as an effective tool to describe and analyse cooperation and competition between political actors. For example, one can represent a Parliament as a signed network of members of parliament (MPs) linked by positive or negative edges, where the sign represents the similarity of their voting behaviours. This can be used to study the dynamics of polarisation among political elites. Such data has been studied in the context of the EU, US, British and Brazilian Parliaments, as well as the United Nations General Assembly [arinikMultiplePartitioningMultiplex2020]; publicly available datasets include signed networks of the EU Parliament [arinikMultiplePartitioningMultiplex2020] and the UK House of Commons [intalDissentRebellionHouse2021]. An approach based on bill co-sponsorship in the US Congress has also been proposed [nealSignTimesWeak2020]. The authors released a collection of signed networks, one for each US congressional session from 1973 (Senate) and 1979 (House of Representatives) up to 2016. The nodes represent MPs, and the sign of the edges is inferred using the SDSM method [nealBackboneBipartiteProjections2014] described in Section 6.1.2. Another interesting data source includes a weighted signed network describing the self-reported similarity between Slovene political parties in 1994 [Kropivnik1996].
@@ 6.2.4 Country systems
International relations studies seek to understand patterns of cooperation, conflict, and interdependence between states. Signed networks provide a natural framework for this, where nodes represent states and signed edges represent cooperative or antagonistic relationships, such as formal alliances, defence agreements, or military disputes [askarisichani19952018GlobalEvolution2020, maozWhatEnemyMy2007, diaz-diazMathematicalModelingLocal2024]. Networks of international relations have also been used as benchmarks to evaluate the performance of balance indices and community detection algorithms [doreianStructuralBalanceSigned2015, estradaWalkbasedMeasureBalance2014, galloTestingStructuralBalance2024, jiangStochasticBlockModel2015, kirkleyBalanceSignedNetworks2019].
@@
  • The Militarised Interstate Disputes dataset (v5.0) records instances in which one state threatened, displayed, or used force against another, covering the period 1816–2014 [palmerMID5Dataset201120142022].
  • The War Data series categorize inter-, intra-, extra-, and non-state wars between 1816–2007 [sarkeesResortWar181620072025].
  • The Defense Cooperation Agreement Dataset (DCAD) [kinne2020defense] covers bilateral defence treaties between 1980–2010, documenting defence-related collaboration such as policy coordination, joint military exercises, peacekeeping operations, or joint research and procurement.
  • The Formal Alliances dataset (v4.1) [giblerInternationalMilitaryAlliances2008] identifies all formal alliances between at least two states between 1816 and 2012, classifying them into defence pacts, neutrality or non-aggression treaties, and entente agreements. This dataset can be used to build positive edges between countries.
Another approach to mapping international relations relies on automated event data collection. The Integrated Crisis Early Warning System (ICEWS) [icews] compiled millions of coded socio-political interactions between international actors, extracted from news reports. ICEWS was discontinued in 2023 and succeeded by the POLitical Event Classification, Attributes, and Types (POLECAT) dataset [polecat], which employs a machine-learning-based coder. POLECAT provides continuously updated, multilingual, and geolocated records of political events from 2018 onward.
Finally, another line of work infers signed relations between countries from cultural, rather than political, interactions [garciaMeasuringCulturalDynamics2013]. One example uses Eurovision voting data to measure affinity or antagonism between countries. The authors define a Friend-or-Foe (FoF) coefficient from voting patterns and compare it with cultural distances computed using Hofstede’s method [hofstede1984culture], finding a significant negative correlation: country pairs with stronger affinities in Eurovision tend to be culturally closer.
@@ 7 Myths and pitfalls
Signed network science is a highly interdisciplinary field, with ideas shared across sociology, physics, computer science, neuroscience, and many other domains. This interdisciplinarity is one of the field’s greatest strengths, but it also creates challenges: terminology can vary across disciplines, concepts sometimes get extended in ways that obscure their original meaning, and the same words may refer to different ideas in different contexts. In this chapter, we discuss claims and assumptions (myths and pitfalls) that we believe benefit from clarification. These are loosely grouped into four themes. Items 1–4 concern the function and interpretation of negative edges. Items 5–8 address the relationship between signed networks and other network types, such as directed, weighted, or multilayer graphs. Items 9–13 deal with claims related to cognitive and structural balance. Finally, items 14–15 focus on other structural aspects. For each, we state the claim, offer our assessment and discuss why the claim may be problematic, where ambiguity arises, and in which cases, if any, it may hold.
@@
    Usually false. In many complex systems, positive and negative links display different structural behaviours [szellMeasuringSocialDynamics2010], have distinct functional properties [fox_PNAS_2005], and are driven by different dynamical mechanisms (see Chapter 5). The sign is therefore indispensable. Ignoring it removes essential information about inhibition, competition, or conflict, leading to misleading interpretations of the system’s organisation and function. As shown in Section 2.1, the coupling between positive and negative subnetworks cannot be treated independently, and as detailed in Section 2.2, many network metrics (clustering, centralities, distances) fundamentally change their interpretation when signs are introduced. While in some contexts focusing on topology alone may be justified, in most applications the sign is a defining element of network structure, behaviour, and function.
@@
    False. The resulting matrix no longer represents the same network. The transformation creates spurious edges between unconnected nodes and treats negative edges as the absence of edges. As a result, the network’s properties become distorted: degree distributions change (Section 2.2), mesoscale organisation is altered (Section 4.4), and dynamical processes defined on the network (Chapter 5) behave fundamentally differently. If a nonnegative representation is required (for example, to apply unsigned spectral methods as in Section 4.4.2), one should instead use mathematically consistent transformations. One example is the Gremban expansion [gremban1996combinatorial, diazdiazGremban], which transforms a signed graph into an unsigned two-layer graph. This method duplicates the node set into two identical layers containing the positive edges, and represents each negative edge as a cross-layer link between the corresponding node copies. This transformation preserves structural properties like balance, mesoscale structure, and walk statistics, while making possible the use of techniques designed for unsigned networks.
@@
    Usually false. Thresholding a signed correlation matrix can be done in two ways, and both approaches introduce systematic distortions. One option is to remove all values below a given threshold without considering their sign, which effectively discards all negative correlations. From the perspective of signed networks, this is particularly problematic: it eliminates all antagonistic or inhibitory relationships that are essential to the system’s structure and dynamics. For instance, in the brain, functional connectivity dynamically partitions into bipartite community structures that coalesce and dissolve over time [sporns2021dynamic]; thresholding away negative correlations would eliminate the anticorrelations that define these bipartitions. The second option is to threshold correlations by absolute value, discarding links of small magnitude but keeping negative values if they are sufficiently strong. While better than erasing all negative links, this approach also presents several issues: it can remove weak but important interactions and systematically bias the inferred network toward denser and more clustered configurations [zalesky2012use]. As a result, thresholding-based network reconstruction is discouraged, since the resulting topology rarely reflects the true organisation of the system [peixoto_PRX_2025].
@@
    False. Signs and directions capture independent but complementary structural dimensions. Signs encode valences (positive or negative), while directions encode asymmetries (who acts on whom). A signed edge does not imply orientation, and a directed edge does not imply positivity or negativity. A possible source of confusion is the different interpretation of negative entries in different matrix representations (Section 2.1). When representing a network through an adjacency matrix, negative entries represent negative edges and asymmetric entries ($A_{ij}\neq A_{ji}$) represent directional links. However, when representing the network through an incidence matrix [biggsAlgebraicGraphTheory1993], positive and negative entries are used to represent the orientation of the edges (+1+1 for head, 1-1 for tail), not their sign. Although most developments in signed network theory have focused on undirected graphs, directionality is present in many real systems. In geopolitical contexts, for instance, both an armed conflict and an invasion may be represented as negative links, yet the direction of interaction carries very different meanings. Directionality can be essential for modelling the dynamical evolution of a network or for interpreting its structural properties. It is not the same for a node to send many negative links as to receive them: in the former case, the node may be conflictive or aggressive, whereas in the latter, it may simply be unpopular. Confounding senders and receivers may thus lead to incorrect conclusions about the network’s functional and dynamical behaviour.
@@
    Misleading. While correlation matrices encode pairwise signed relationships, they do not automatically constitute networks with meaningful structure. That said, network-based methods can be used to extract modules from a correlation matrix [macmahonCommunityDetectionCorrelation2015], interpreting correlations as positive/negative network links. This approach is, however, only methodological: it employs a network representation to facilitate the analysis of the data, but does not imply that the underlying system is itself a network. Note, however, that it is possible to infer a signed network from a correlation matrix, but only after applying appropriate statistical methods such as Bayesian inference [peixoto_PRX_2025].
@@
    False. Structural balance describes a mathematical property of network organisation, not the quality of social relations or the absence of conflict. A perfectly balanced network naturally organises into two (or more, under weak balance) mutually antagonistic factions with exclusively positive edges within groups and negative edges between them (Section 3.4). Empirical studies of real social networks show that balanced states often correspond to the most polarised or fragmented configurations [talagaPolarizationMultiscaleStructural2023, diaz-diazMathematicalModelingLocal2024, vendevilleModelingEchoChamber2025]. Conversely, networks with low or moderate balance may exhibit more blended dynamics where positive and negative relationships coexist in less extreme patterns, potentially indicating more complex but less polarised social structures [talagaPolarizationMultiscaleStructural2023]. The apparent paradox—that ”balance” produces polarisation—arises from conflating the colloquial meaning of balance (harmony, moderation) with its precise structural definition (even numbers of negative edges in cycles). As discussed in Section 3.3, this confusion partly comes from the different perspectives of Heiderian cognitive balance (which does emphasise psychological consistency and reduced tension at the individual level) versus Hararian structural balance (which characterises global network organisation that may produce group-level antagonism). Interpreting structural balance as evidence of peaceful coexistence or social harmony is misleading, and we should instead recognise it as a signature of structural division.
@@
    Only partially true. Structural balance measures how easily a network can be divided into antagonistic groups with positive internal and negative external links, and a network is close to balance when only a few edges or cycles violate this pattern. This descriptive view does not require comparison with a null model and expresses how far from perfect balance a signed network is. However, null models become necessary when the goal is to assess significance rather than describing the network. They indicate whether the observed balance goes beyond what a random sign assignment would produce, and thus whether the pattern reflects genuine structure or mere noise. The design of null models must also match the system: for example, if links can be created and destroyed, both topology and signs should be randomised, whereas if connections are fixed, only signs should be (see Section 4.1). In short, both views are complementary: descriptive measures quantify how balanced a network is, while null models test whether this level of balance is meaningful. This distinction mirrors the descriptive versus inferential approaches to community detection [peixotoDescriptiveVsInferential2023].
@@
    Usually false. This misconception operates at two levels: conceptual (what communities are) and methodological (how to find them). Conceptually, most signed network literature defines communities as groups with predominantly positive edges within and negative edges between them, following principles of structural balance (Section 4.4). This differs fundamentally from unsigned networks, where communities are simply densely connected subgraphs without considering edge signs. However, some works apply unsigned definitions to signed data, leading to confusion when methods correctly identify signed communities (factions) but are dismissed as incorrect [knyazevSpectralPartitioningSigned2018, foxInvestigationSpectralClustering2018]. A terminological proposal [diazdiazGremban] helps clarify: use ”communities” for densely connected groups regardless of sign, and ”factions” for groups with positive internal and negative external edges. This separation makes explicit that dense connectivity and sign-based cohesion are distinct structural properties. Methodologically, no single algorithm is universally correct for detecting mesoscale structure in signed networks, just as in unsigned networks. The appropriate method depends on the goal of the analysis and the assumptions about the data. For example, spectral methods are well suited when the aim is to find factions that follow structural balance and with approximately the same size. In contrast, stochastic block models are more flexible and can detect a wider range of patterns, including signed core–periphery structures and other configurations that depart from strict balance.

### v2

@@ 6.2.1 In-person social communities
  • Sampson’s monastery data [339] records the evolution of relations between 18 monks during a political crisis within a US cloister, across three temporal waves. Each wave, the monks were asked to select the three others they esteemed the most and the one they esteemed the least. Similar ratings were collected for liking, influence, and praise instead of esteem.
  • Nordlie-Newcomb’s fraternity dataset [298, 292] is the result of an experiment of social psychology conducted at the University of Michigan. Seventeen male college students were housed together, and the authors periodically collected data via surveys to assess the evolution of friendships and antagonisms between the students. A similar study collected data about college female students in the US [228].
  • Spanish high school students [332]. Recently, Ruiz-Garcia et al. [332] surveyed more than three thousand students (11-15 years old) in 13 Spanish schools, asking them about which other students they liked or disliked. The result is a directed, weighted network with over 60,000 edges, with additional information about individual features and personality traits.
  • Behavioural observations on preschoolers [286]. Neal et al. [286] inferred positive and negative links between children in two classrooms of preschoolers ($N=35$), based on behavioural patterns during playtime. To do so, they leverage the backbone procedure described in [291] (see Section 6.1.2).
Other relevant data sources include signed networks of relationships between antique Philosophers [228], alliances and antagonisms between $N=16$ tribes of the Highlands region in New Guinea [324], as well as relationships between inhabitants of $N=32$ Honduran villages collected during a clinical trial. We finally mention another widely used source of data, the Urban Communes Dataset, which records the nature of relationships between groups of inhabitants in several major US cities, alongside socio-demographic indicators and individual beliefs, across multiple waves. Formerly freely available, it is now accessible on request.
@@ 6.2.2 Online platforms
The availability of large-scale, high-quality signed datasets for studying online social interactions remains limited. A few datasets have become standard benchmarks, in particular those hosted in the SNAP repository [233], and have supported a wide range of work on sign prediction, community detection, and structural balance:
  • Epinions trust network [232]. This dataset originates from the Epinions product review platform, where users could explicitly express trust or distrust toward other members. The resulting signed network captures these declared trust relationships among users.
  • Slashdot Zoo 313131www.slashdot.org [148, 232]. Slashdot is a technology-oriented news platform featuring user-submitted and editor-evaluated stories related to science and technology. In 2002, the site introduced the Zoo feature, allowing users to tag others as friends (those whose comments they appreciated) or foes (those whose comments they found uninteresting). The dataset includes three temporal snapshots of this signed social network.
Both Epinions [245, 231, 328, 2, 303] and Slashdot [223, 2, 303] have been widely used in the literature for sign prediction and structural balance studies.
  • Bitcoin Alpha and Bitcoin OTC [222]. These datasets represent who-trusts-whom networks extracted from two Bitcoin trading platforms. Users rate each other following transactions, creating directed weighted links that indicate trust or distrust. Since participants are anonymous, the reputation systems underlying these platforms aim to mitigate fraud and risk.
  • Wikipedia Requests for Adminship (RfA) [231, 232, 396]. This dataset captures the voting process through which Wikipedia editors decide whether to promote a candidate to administrator status. Each vote can express support, opposition, or neutrality. The initial version, based on a 2008 dump, includes nearly 2,8002{,}800 elections. A later extended version by West et al. [396] spans the entire period from 2003 to 2013 and also contains the textual content of votes (i.e., the accompanying comments). Besides RfA, several studies have inferred signed relations from co-editing behaviour and interaction patterns among editors (see Section 6.1.4).
  • Reddit Hyperlink Network [222]. This dataset represents interactions between subreddits over a 2.5-year period (2014–2017). It is constructed from hyperlinks in user posts that reference content in other subreddits. Each edge originates from a source subreddit and points to a target subreddit, with sentiment annotations derived from the text of the linking post, resulting in a signed inter-community network.
@@
  • DEBAGREEMENT [316]. This dataset consists of 42,894 annotated comment–reply pairs collected from five politically and socially active subreddits: r/BlackLivesMatter, r/Brexit, r/climate, r/democrats, and r/Republican. Each pair is labelled as agree, neutral, or disagree, creating a signed interaction network among users based on the tone of their replies.
  • Birdwatch (now Community Notes) [407, 6]. This platform was launched by Twitter in January 2021 as a crowd-sourced fact-checking platform where selected users rated the trustworthiness of Tweets through evaluative notes. Signed networks have been extracted from user-note ratings from Birdwatch data to study political polarisation and antagonism [407, 6]. As of this writing, anonymised Birdwatch (now Community Notes) data remains publicly available.
  • DerStandard Forum Interactions [130, 132]. DerStandard  [^32] https://www.derstandard.at/ is an Austrian online newspaper that features a popular comment section integrated with its articles. The dataset captures user interactions on the platform over nearly ten years (2013–2022), including explicit agreement and disagreement votes on comments. These interactions can be used to build temporal signed networks across diverse topics such as football, national elections, or COVID-19 policy [131].
  • Meneame.net. Meneame  [^33] https://www.meneame.net/, a Spanish news aggregator with a similar structure to Reddit, was studied by [54, 53] to investigate the role of negative interactions in highlighting polarised content. The platform allows users to submit news stories and interact through upvotes, downvotes, and comments. Moreover, unlike many platforms, Menéame applies no personalised recommendations, so any observed phenomenon arises purely from user behaviour.
Some additional platforms have supported signed or sign-inferable interactions, though the corresponding research does not provide public access to the datasets. Examples are from Yahoo Answers [405], where the authors infer trust and distrust through question-and-answer behaviour and voting patterns; Essembly.com [52], where users can label others as “friends” or “foes”; Pardus [370], a multi-relational online game, included both cooperative and antagonistic links.
@@ 6.2.3 Political systems
The formalism of signed networks has raised interest as an effective tool to describe and analyse cooperation and competition between political actors. For example, one can represent a Parliament as a signed network of members of parliament (MPs) linked by positive or negative edges, where the sign represents the similarity of their voting behaviours. This can be used to study the dynamics of polarisation among political elites. Such data has been studied in the context of the EU, US, British, and Brazilian Parliaments, as well as the United Nations General Assembly [24]; publicly available datasets include signed networks of the EU Parliament [24] and the UK House of Commons [197]. An approach based on bill co-sponsorship in the US Congress has also been proposed [289]. The authors released a collection of signed networks, one for each US congressional session from 1973 (Senate) and 1979 (House of Representatives) up to 2016. The nodes represent MPs, and the sign of the edges is inferred using the SDSM method [291] described in Section 6.1.2. Another interesting data source includes a weighted signed network describing the self-reported similarity between Slovene political parties in 1994 [218].
@@ 6.2.4 Country systems
International relations studies seek to understand patterns of cooperation, conflict, and interdependence between states. Signed networks provide a natural framework for this, where nodes represent states and signed edges represent cooperative or antagonistic relationships, such as formal alliances, defence agreements, or military disputes [26, 258, 90]. Networks of international relations have also been used as benchmarks to evaluate the performance of balance indices and community detection algorithms [99, 106, 138, 202, 212].
@@
  • The Militarised Interstate Disputes dataset (v5.0) records instances in which one state threatened, displayed, or used force against another, covering the period 1816–2014 [301].
  • The War Data series categorize inter-, intra-, extra-, and non-state wars between 1816–2007 [340].
  • The Defense Cooperation Agreement Dataset (DCAD) [211] covers bilateral defence treaties between 1980–2010, documenting defence-related collaboration such as policy coordination, joint military exercises, peacekeeping operations, or joint research and procurement.
  • The Formal Alliances dataset (v4.1) [145] identifies all formal alliances between at least two states between 1816 and 2012, classifying them into defence pacts, neutrality or non-aggression treaties, and entente agreements. This dataset can be used to build positive edges between countries.
Another approach to mapping international relations relies on automated event data collection. The Integrated Crisis Early Warning System (ICEWS) [48] compiled millions of coded socio-political interactions between international actors, extracted from news reports. ICEWS was discontinued in 2023 and succeeded by the POLitical Event Classification, Attributes, and Types (POLECAT) dataset [341], which employs a machine-learning-based coder. POLECAT provides continuously updated, multilingual, and geolocated records of political events from 2018 onward.
Finally, another line of work infers signed relations between countries from cultural, rather than political, interactions [142]. One example uses Eurovision voting data to measure affinity or antagonism between countries. The authors define a Friend-or-Foe (FoF) coefficient from voting patterns and compare it with cultural distances computed using Hofstede’s method [180], finding a significant negative correlation: country pairs with stronger affinities in Eurovision tend to be culturally closer.
@@ 7 Myths and pitfalls
Signed network science is a highly interdisciplinary field, with ideas shared across sociology, physics, computer science, neuroscience, and many other domains. This interdisciplinarity is one of the field’s greatest strengths, but it also creates challenges: terminology can vary across disciplines, concepts sometimes get extended in ways that obscure their original meaning, and the same words may refer to different ideas in different contexts. In this chapter, we discuss claims and assumptions (myths and pitfalls) that we believe benefit from clarification. These are loosely grouped into four themes. Items 1–4 concern the function and interpretation of negative edges. Items 5–8 address the relationship between signed networks and other network types, such as directed, weighted, or multilayer graphs. Items 9–13 deal with claims related to cognitive and structural balance. Finally, items 14–15 focus on other structural aspects. For each, we state the claim, offer our assessment, and discuss why the claim may be problematic, where ambiguity arises, and in which cases, if any, it may hold.
@@
    Usually false. In many complex systems, positive and negative links display different structural behaviours [371], have distinct functional properties [129], and are driven by different dynamical mechanisms (see Chapter 5). The sign is therefore indispensable. Ignoring it removes essential information about inhibition, competition, or conflict, leading to misleading interpretations of the system’s organisation and function. As shown in Section 2.1, the coupling between positive and negative subnetworks cannot be treated independently, and as detailed in Section 2.2, many network metrics (clustering, centralities, distances) fundamentally change their interpretation when signs are introduced. While in some contexts focusing on topology alone may be justified, in most applications the sign is a defining element of network structure, behaviour, and function.
@@
    False. The resulting matrix no longer represents the same network. The transformation creates spurious edges between unconnected nodes and treats negative edges as the absence of edges. As a result, the network’s properties become distorted: degree distributions change (Section 2.2), mesoscale organisation is altered (Section 4.4), and dynamical processes defined on the network (Chapter 5) behave fundamentally differently. If a nonnegative representation is required (for example, to apply unsigned spectral methods as in Section 4.4.2), one should instead use mathematically consistent transformations. One example is the Gremban expansion [154, 91], which transforms a signed graph into an unsigned two-layer graph. This method duplicates the node set into two identical layers containing the positive edges, and represents each negative edge as a cross-layer link between the corresponding node copies. This transformation preserves structural properties like balance, mesoscale structure, and walk statistics, while making possible the use of techniques designed for unsigned networks.
@@
    Usually false. Thresholding a signed correlation matrix can be done in two ways, and both approaches introduce systematic distortions. One option is to remove all values below a given threshold without considering their sign, which effectively discards all negative correlations. From the perspective of signed networks, this is particularly problematic: it eliminates all antagonistic or inhibitory relationships that are essential to the system’s structure and dynamics. For instance, in the brain, functional connectivity dynamically partitions into bipartite community structures that coalesce and dissolve over time [363]; thresholding away negative correlations would eliminate the anticorrelations that define these bipartitions. The second option is to threshold correlations by absolute value, discarding links of small magnitude but keeping negative values if they are sufficiently strong. While better than erasing all negative links, this approach also presents several issues: it can remove weak but important interactions and systematically bias the inferred network toward denser and more clustered configurations [411]. As a result, thresholding-based network reconstruction is discouraged, since the resulting topology rarely reflects the true organisation of the system [311].
@@
    False. Signs and directions capture independent but complementary structural dimensions. Signs encode valences (positive or negative), while directions encode asymmetries (who acts on whom). A signed edge does not imply orientation, and a directed edge does not imply positivity or negativity. A possible source of confusion is the different interpretation of negative entries in different matrix representations (Section 2.1). When representing a network through an adjacency matrix, negative entries represent negative edges and asymmetric entries ($A_{ij}\neq A_{ji}$) represent directional links. However, when representing the network through an incidence matrix [38], positive and negative entries are used to represent the orientation of the edges (+1+1 for head, 1-1 for tail), not their sign. Although most developments in signed network theory have focused on undirected graphs, directionality is present in many real systems. In geopolitical contexts, for instance, both an armed conflict and an invasion may be represented as negative links, yet the direction of interaction carries very different meanings. Directionality can be essential for modelling the dynamical evolution of a network or for interpreting its structural properties. It is not the same for a node to send many negative links as to receive them: in the former case, the node may be conflictive or aggressive, whereas in the latter, it may simply be unpopular. Confounding senders and receivers may thus lead to incorrect conclusions about the network’s functional and dynamical behaviour.
@@
    Misleading. While correlation matrices encode pairwise signed relationships, they do not automatically constitute networks with meaningful structure. That said, network-based methods can be used to extract modules from a correlation matrix [255], interpreting correlations as positive/negative network links. This approach is, however, only methodological: it employs a network representation to facilitate the analysis of the data, but does not imply that the underlying system is itself a network. Note, however, that it is possible to infer a signed network from a correlation matrix, but only after applying appropriate statistical methods such as Bayesian inference [311].
@@
    False. Structural balance describes a mathematical property of network organisation, not the quality of social relations or the absence of conflict. A perfectly balanced network naturally organises into two (or more, under weak balance) mutually antagonistic factions with exclusively positive edges within groups and negative edges between them (Section 3.4). Empirical studies of real social networks show that balanced states often correspond to the most polarised or fragmented configurations [374, 90, 391]. Conversely, networks with low or moderate balance may exhibit more blended dynamics where positive and negative relationships coexist in less extreme patterns, potentially indicating more complex but less polarised social structures [374]. The apparent paradox—that ”balance” produces polarisation—arises from conflating the colloquial meaning of balance (harmony, moderation) with its precise structural definition (even numbers of negative edges in cycles). As discussed in Section 3.3, this confusion partly comes from the different perspectives of Heiderian cognitive balance (which does emphasise psychological consistency and reduced tension at the individual level) versus Hararian structural balance (which characterises global network organisation that may produce group-level antagonism). Interpreting structural balance as evidence of peaceful coexistence or social harmony is misleading, and we should instead recognise it as a signature of structural division.
@@
    Only partially true. Structural balance measures how easily a network can be divided into antagonistic groups with positive internal and negative external links, and a network is close to balance when only a few edges or cycles violate this pattern. This descriptive view does not require comparison with a null model and expresses how far from perfect balance a signed network is. However, null models become necessary when the goal is to assess significance rather than describing the network. They indicate whether the observed balance goes beyond what a random sign assignment would produce, and thus whether the pattern reflects genuine structure or mere noise. The design of null models must also match the system: for example, if links can be created and destroyed, both topology and signs should be randomised, whereas if connections are fixed, only signs should be (see Section 4.1). In short, both views are complementary: descriptive measures quantify how balanced a network is, while null models test whether this level of balance is meaningful. This distinction mirrors the descriptive versus inferential approaches to community detection [310].
@@
    Usually false. This misconception operates at two levels: conceptual (what communities are) and methodological (how to find them). Conceptually, most signed network literature defines communities as groups with predominantly positive edges within and negative edges between them, following principles of structural balance (Section 4.4). This differs fundamentally from unsigned networks, where communities are simply densely connected subgraphs without considering edge signs. However, some works apply unsigned definitions to signed data, leading to confusion when methods correctly identify signed communities (factions) but are dismissed as incorrect [216, 128]. A terminological proposal [91] helps clarify: use ”communities” for densely connected groups regardless of sign, and ”factions” for groups with positive internal and negative external edges. This separation makes explicit that dense connectivity and sign-based cohesion are distinct structural properties. Methodologically, no single algorithm is universally correct for detecting mesoscale structure in signed networks, just as in unsigned networks. The appropriate method depends on the goal of the analysis and the assumptions about the data. For example, spectral methods are well-suited when the aim is to find factions that follow structural balance and are approximately the same size. In contrast, stochastic block models are more flexible and can detect a wider range of patterns, including signed core–periphery structures and other configurations that depart from strict balance.